OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\dfrac{a}{b}=\dfrac{b}{c}=\dfrac{c}{d} \hspace{.33em}\\~\\ \normalsize \text{then} \hspace{.33em}\\~\\ \dfrac{a-d}{b-c}\geq x \hspace{.33em}\\~\\ \normalsize \text{find x} \hspace{.33em}\\~\\ a.)2 \hspace{.33em}\\~\\ b.)3 \hspace{.33em}\\~\\ c.)1 \hspace{.33em}\\~\\ d.)0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\Large\tt \begin{align} \color{black}{\dfrac{a}{b}=\dfrac{b}{c}=\dfrac{c}{d}=k \hspace{.33em}\\~\\~\\ \underline{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~}\\~\\ \dfrac{a-d}{b-c} \hspace{.33em}\\~\\ =\dfrac{bk-\frac{c}{k}}{ck-c} \hspace{.33em}\\~\\ =\dfrac{(ck)k-\frac{c}{k}}{ck-c} \hspace{.33em}\\~\\ =\dfrac{k^2-\frac{1}{k}}{k-1} \hspace{.33em}\\~\\ =\dfrac{k^3-1}{k(k-1)} \hspace{.33em}\\~\\ =\dfrac{k^3-1}{k(k-1)} \hspace{.33em}\\~\\ =\dfrac{(k-1)(k^2+k+1)}{k(k-1)} \hspace{.33em}\\~\\ =1+k+\dfrac{1}{k} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{f(x)=1+x+\dfrac{1}{x} \hspace{.33em}\\~\\ f(x)'=1-\dfrac{1}{x^2} \hspace{.33em}\\~\\ 0=1-\dfrac{1}{x^2} \hspace{.33em}\\~\\ x=\pm1 \hspace{.33em}\\~\\ f(x)_{max}=1+1+\dfrac{1}{1} \hspace{.33em}\\~\\ f(x)_{max}=3 }\end{align}\)

OpenStudy (kainui):

A slightly different way I did it was this: \[\Large a=k^3d \\ \Large b=k^2d \\ \Large c=kd\] Plug them in \[\Large \frac{k^3d-d}{k^2d-kd}=\frac{k^3-1}{k^2-k}= \frac{k^2+k+1}{k}=k+1+\frac{1}{k}\] Then you can finish just lie you did, or take the cool, completely ridiculous route I took: \[\Large 1+k+\frac{1}{k} = 1+2 \cosh (\ln(k))\] Then take the derivative of this \[\Large \frac{d}{dk} \left( 1+2 \cosh (\ln(k)) \right) =2 \sinh(\ln(k))\frac{1}{k}=0\] So either sinh or 1/k is 0, but we obviously don't want infinity, so we throw away the 1/k part (along with dividing by 2) \[\Large \sinh(\ln(k))=0\] Similar to sin(0)=0, sinh(0)=0 so ln(k) must be 0 as well. \[\Large \ln(k)=0\] And of course, log of 1 is 0, so k=1, just like you got. Just a weird way to the same answer.

OpenStudy (mathmath333):

looks cool i dont get why u took \(\large \begin{align} \color{black}{ 1+k+\dfrac{1}{k} = 1+2 \cosh (\ln(k)) \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (kainui):

I just wanted to do something weird, the main trick I wanted to show you was how I made everything k^p*d and plugged it in at the start, it made the calculation much cleaner. That second part I was just bored messing around and thought it would be interesting to see. =P

OpenStudy (mathmath333):

ok , i prefer any method which is short and lazy